# ”They’ve Stolen My GPL-Licensed Model!”: Toward Standardized and Transparent Model Licensing

### v1

@@ Abstract.
As model parameter sizes reach the billion-level range and their training consumes zettaFLOPs of computation, components reuse and collaborative development are become increasingly prevalent in the Machine Learning (ML) community. These components, including models, software, and datasets, may originate from various sources and be published under different licenses, which govern the use and distribution of licensed works and their derivatives. However, commonly chosen licenses, such as GPL and Apache, are software-specific and are not clearly defined or bounded in the context of model publishing. Meanwhile, the reused components may also have free-content licenses and model licenses, which pose a potential risk of license noncompliance and rights infringement within the model production workflow. In this paper, we propose addressing the above challenges along two lines: 1) For license analysis, we have developed a new vocabulary for ML workflow management and encoded license rules to enable ontological reasoning for analyzing rights granting and compliance issues. 2) For standardized model publishing, we have drafted a set of model licenses that provide flexible options to meet the diverse needs of model publishing. Our analysis tool is built on Turtle language and Notation3 reasoning engine, envisioned as a first step toward Linked Open Model Production Data. We have also encoded our proposed model licenses into rules and demonstrated the effects of GPL and other commonly used licenses in model publishing, along with the flexibility advantages of our licenses, through comparisons and experiments.
License Analysis, AI Licensing, Automated Reasoning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Social and professional topics Testing, certification and licensingccs: Software and its engineering Open source modelccs: Theory of computation Automated reasoning
@@ 1. Introduction
In recent years, the compelling generalization capabilities provided by billion-parameter models (Wei et al., 2023), along with the high computational and data costs associated with their training (Maslej et al., 2024), have motivated ML project developers to collaborate incrementally rather than train models from scratch. For example, a common approach is to download a Pre-Trained Model (PTM) (Jiang et al., 2023) and fine-tune it for downstream task (Hu et al., 2022). However, these paradigms may face potential legal risks if the use and redistribution practices violate the governing licenses of the reused components, akin to the GPL violation issues in the field of Open Source Software (OSS) (Mathur et al., 2012). Another risk arises from the choice of license used to republish the work. Some developers adhere to traditional software publishing practices and select OSS licenses for their models (Devlin et al., 2019; Ni et al., 2022), which often lack clear definitions and conditions regarding ML activities and do not effectively prevent undesirable use. For example, a licensee can close-source your published models, even if they are licensed under GPL, without violating any terms.
@@
There are three possible ways to mitigate above risks. First, developers could avoid using any third-party materials. However, this is extremely difficult for individual developers, as training PTMs is expensive and requires vast amounts of data. For instance, the training dataset for GPT-2 (Radford et al., 2019) was collected from 45 million web pages, governed by various licenses and terms of use. Second, a new publishing standard for ML projects could be developed, which might include drafting specific licenses for models and datasets (Benjamin et al., 2019; Contractor et al., 2022), along with a compatibility table to guide their reuse policies. However, this approach also has several limitations, as it does little to address existing conflicts in ML projects that already rely on components released under traditional licenses. Furthermore, it is impractical to expect all publishers to relicense their previous works. Third, we can scan the reused components in ML projects and analyze existing license compliance risks to eliminate them. This is a common solution applied to OSS projects (Jaeger et al., 2017) but it cannot be directly extended to ML projects. The reason is that ML components can involve complex coupling mechanisms and different licensing frameworks that are interwoven within a project.
@@
Take MixLoRA (Li et al., 2024) as an example: it is licensed under Apache-2.0 (an OSS license) and is fine-tuned on Llama 2 model (Touvron et al., 2023) (governed by Llama 2 Community License (Meta Platforms, 2024), a model license) using the Cleaned Alpaca Dataset (Taori et al., 2023), which licensed under CC BY-NC-4.0 (a free-content license from Creative Commons, aka CC (Commons, 2024)). Previous OSS license analysis tools (Mathur et al., 2012; Ombredanne, 2020) that only consider package reference dependencies and focus on software licenses will fall short in such ML scenarios. Therefore, to provide license analysis for ML projects, the key is to develop a specific ontology that describes the ML workflow and provides a corresponding interpretative solution for licenses, covering all licensing frameworks and disambiguating their mapping rules related to ML activities. Moreover, the lack of consensus in standard model publishing practices and the inflexibility of existing publicly available model licenses have led many developers to publish their models under OSS licenses or even free-content licenses (Groeneveld et al., 2024; CohereForAI, 2024), further complicating the design of license analysis methods.
@@
In this paper, we propose a two-pronged approach to address these challenges. First, to resolve existing noncompliance in ML projects, we introduce MG Analyzer, a tool that constructs ML workflows as Resource Description Framework (RDF) (Pan, 2009) graphs and assesses potential license compliance issues, improper license selection, granting of rights, restrictions, and obligations within the projects. Second, to promote standardized model publishing in the future, we propose a new set of model-specific licenses, MG Licenses, offering CC-style licensing options for developers to choose from. To present potential risks of using traditional OSS, model, and free-content licenses in model publishing scenarios, we evaluate them with the MG Analyzer on a typical workflow. We also demonstrate the flexibility of MG Licenses in encompassing nearly all licensing conditions provided by other model licenses through comprehensive comparisons.
The main contributions of our paper are:
@@
  • We identified the challenges of license compliance and model licensing in ML projects.
@@
  • We developed MG Analyzer using semantic technologies to automate license analysis in ML projects. It includes a vocabulary for describing ML workflows with dependencies and license rules. We also provide a interface to convert user-input workflow descriptions into RDF graphs following this vocabulary. Users can convert workflow descriptions into RDF graphs, enabling the tool to construct dependencies, reason, and detect license conflicts.
@@
  • To the best of our knowledge, MG Analyzer and Licenses represent the first attempt at standardizing model publishing. The proposed code and license drafts are available at this link.
The rest of the paper is organized as follows. Section 2 presents background and related studies. Section 3 introduces our proposed vocabulary for workflow descriptions and the license analysis tool. Section 4 offers a comprehensive comparison of commonly used licenses and briefly outlines the advantages of the new model licenses we propose. Section 5 presents license analysis results to demonstrate the risks associated with non-standard licensing, while Section 6 concludes the paper. Supplementary tables and codes are provided in the Appendix.
@@ 2. Background and Related Work
2.1. License Compliance Analysis
The previous license compliance analysis studies primarily focus on OSS-licensed software (Schoettle, 2019; Cui et al., 2023), and several successful tools, such as FOSSology (Jaeger et al., 2017) and Black Duck Software Composition Analysis (Inc., 2024), have been developed. The main goal of these tools is to identify all open-source dependencies in software projects to evaluate associated license compliance risks, obligations, and attribution requirements. Typically, the component dependencies and license information in a software project can be obtained through scanning and matching (Ombredanne, 2020). This process involves gathering information from notices, headers, licenses, and other project files or attempting to match the code to determine its provenance. Unfortunately, these strategies cannot be naturally extended to ML projects for the following reasons.
First, the components of ML projects, particularly models, have more intricate dependencies than code. For example, knowledge can be transferred between models without explicitly copying weights (You et al., 2021). Second, the OSS Bill of Materials standard, such as Software Package Data Exchange (SPDX) (Foundation, 2024), does not fully support common model licenses like OpenRAIL-M (Contractor et al., 2022), Llama 2 (Meta Platforms, 2024). Third, non-standard licensing (model publishing under OSS or free-content licenses, refer to the statistical data in Table 5) is prevalent in ML projects (Duan et al., 2024), adding complexity to license analysis. Furthermore, the crowd-sourced nature of ML components may also lead to over-permissive licenses (Rajbahadur et al., 2021), distorting the analysis results.
For these reasons, license compliance issues in the ML field remain nearly unexplored. Rajbahadur et al. (Rajbahadur et al., 2021) investigated license provenance issues in ML datasets and found instances of non-compliance between their licenses and the licenses of their data sources. Building on these findings, Duan et al. (Duan et al., 2024) proposed a tool that analyzes conflicts directly based on the licenses of data sources and provided guidelines to minimize such conflicts. However, their tool is limited to generating analysis reports and lacks the ability to visualize and exchange ML workflows, making it difficult to extend or integrate external resources from the web (e.g., linking to another workflow via URI). Therefore, in this work, we propose an ontology describing ML workflows using RDF graphs (a visualized workflow is provided in Appendix E), making it both extensible and linkable, and then analyze license compliance through an automated reasoning engine.
2.2. Model Licensing
@@ 3. MG Analyzer
Figure 1. Overview of MG Analyzer. (”mg” is the prefix for our proposed vocabulary.)
[…]
This section aims to introduce the specific design of MG Analyzer by exploring three questions: (i) How can we represent the workflows of ML projects? (ii) How do we establish a mapping from license text to reasonable rules? (iii) What types of license compliance issues can arise in ML projects, and how can we detect them? Before delving into the detailed design that answers these questions, we first provide an overview that serves as a roadmap for this section. As illustrated in Figure 1, the process of MG Analyzer is divided into three main parts: Construction, Reasoning, and Analysis.
@@ 3.1. ML Workflow Representation
① ML workflows often involve various components (e.g., code, datasets, images, model weights, services), each governed by licenses from different frameworks. Additionally, non-standard licensing practices are prevalent in current ML projects (Rajbahadur et al., 2021; Duan et al., 2024), for instance, the C4AI Command R+ model (CohereForAI, 2024) is licensed under a free-content license: CC-BY-NC-4.0. Therefore, the representation should be flexible enough to cover such situations.
@@
At this stage, we can describe a base ML workflow, as illustrated111 The ”mg” prefix is omitted and some properties are merged or filtered for presentation. schematically in Figure 2 (The RDF graph can be referred to in gen.ttl in Figure 1). In this base workflow, the derived input and output of each Action can be represented as blank nodes, serving as placeholders. These placeholders will be populated by reasoning the output yielded by the previous action and the input for the current action, respectively.
@@
The rights-using dependencies describe the rights that must be granted for actions performed on works. For example, when executing a Train action on a model, it requires the rights to use and modify (termed as Usage in MG Vocabulary) from the model’s license222 An example of such rights granting can be found in the Llama 2 Community License, which states, ”You are granted a … to use, reproduce, distribute, copy, create derivative works of, and make modifications to the Llama Materials.”.. Similarly, the rights-using dependencies should proliferate according to compositional dependencies. For instance, the requirement to modify a model extends to all its submodels. In the MG Analyzer, we create new nodes called Request to represent this dependency. Additionally, it is insufficient to only check the granting rights; the reserved rights must also be verified. Depending on the clarity of the license text, some rights may either be explicitly granted or reserved. Furthermore, certain license clauses can waive the requirement for specific rights. For instance, both GPL-3.0 and CC licenses include automatic relicensing clauses for downstream recipients, which eliminate the need for a sublicense right.
@@ 3.2. License Rule Encoding and Reasoning
Table 1. Supported Actions in MG Analyzer Following Rule Alignment. The symbols $=$ and $\approx$ indicate that the Type/Form of output work and input work are the same and may differ, respectively. The corresponding OSS, Free-content, and Model license terms for each action are listed.
Action Type Form Composition Terms
Copy $=$ $=$ Output and input are exactly same. Copy Duplicate
Modify $=$ $=$ Output includes a significant portion of input and can be reverted. Modify Fine-tune
Amalgamate $=$ $=$ Output includes portions of input but cannot be reverted. Modify Remix Fusion
Train $=$ $=$ Output has the same structure as input and may contain a negligible portion of it. Alter Adapt Train
Synthetic
Distill $=$ $=$ Output does not contain any portion of input but performs similarly to it. Distill Transfer
Extract
Embed $=$ $=$ Output does not contain any portion of input, but there has a mapping that converts input to output. Translate Transform
Publish $=$ $\approx$ Output is the same as the input but may have a different form. Redistribute Perform Display Disseminate
@@
The final aspect is applicable term generalization, where we encode the triggering conditions of a license term into the following properties: range of input work forms, range of output work forms, and types of actions. Figure 3 illustrates an example of the GPL-3.0 derivatives rule333 The original GPL-3.0 license text reads: ”You may convey a work based on the Program … in the form of source code … provided that you also meet all of these conditions: … stating that you modified it … it is released under this License … keep intact all notices … license the entire work, as a whole, under this License …” ., which represents the Combine action applied to input works in code format, resulting in a code output. This action triggers the ”derivative” clause in GPL-3.0, indicating that the license of the output work must be compatible with GPL-3.0 (e.g., APGL-3.0). Additionally, five restrictions must apply to the output work if it is to be republished, as dictated by this definition dependency. This rule does not include any Use Restrictions, which apply to output works regardless of whether they are republished. We found this subtle distinction to be crucial in ML license analysis, as most OSS license terms are triggered by distribution, and their definitions of ”distribution” typically exclude publishing as a service. However, in the case of models, the common deployment method is through a web interface, leading to many OSS license restrictions being circumvented in such scenarios (ref Section 5).
Furthermore, multiple rules may lead to the same output definition, and we provide the option of fuzz form matching by enabling fuzz rules to enhance the interpretive capabilities of the MG Analyzer. It is worth mentioning that the reasoning results behind the restrictions derived from these rules require further analysis for validation. Taking the rule in Figure 3 as an example, the presence of additional discrimination terms in the final work that violate GNU freedom (Greenbaum, 2016) dictates whether errors related to GNU freedom conflicts should be reported. We present snippets of our encoded rules written in Turtle format in Appendix C. The list of supported licenses, whose terms have been encoded in MG Analyzer, is shown in Table 3, covering nearly all top-ranking licenses for published models on HuggingFace444 https://huggingface.co/models, a popular model publishing platform.
@@
[…]
With MG Vocabulary, the ML workflow with compositional dependencies, and encoded license rules, we can reason to derive the definition and rights-using dependencies, thereby determining the applicable licenses for intermediate works. The license determination in the MG Analyzer follows an incremental and minimal noncompliance strategy, where the new license only applies to the incremental parts of the work without affecting the original work (a common practice in licensing). Furthermore, to avoid introducing additional compliance issues during analysis, we use Unlicense as the default license when applicable. However, an exception arises with the license proliferation clauses found in copyleft licenses, such as GPL-3.0 and OSL-3.0, which require that the entire new work be licensed under the same terms. Our analyzer incorporates reasoning logic to identify applicable licensing solutions (a snippet of logic can be found in Appendix C), but unresolved conflicts may occur if multiple copyleft clauses are triggered. In such cases, the MG Analyzer will select one of these copyleft licenses and report an error during the Analysis Stage.

### v2

@@ Abstract.
As model parameter sizes scale into the billions and training consumes zettaFLOPs of computation, the reuse of Machine Learning (ML) assets and collaborative development have become increasingly prevalent in the ML community. These ML assets, including models, datasets, and software, may originate from various sources and be published under different licenses, which govern the use and distribution of licensed works and their derivatives. However, commonly chosen licenses, such as GPL and Apache, are software-specific and are not clearly defined or bounded in the context of model publishing. Meanwhile, the reused assets may also be under free-content licenses and model licenses, which pose a potential risk of license noncompliance and rights infringement within the model production workflow. In this paper, we address these challenges along two lines: 1) For ML workflow compliance, we propose ModelGo (MG) Analyzer, a tool that incorporates a vocabulary for ML workflow management and encoded license rules, enabling ontological reasoning to analyze rights granting and compliance issues. 2) For standardized model publishing, we introduce ModelGo Licenses, a set of modell-specific licenses that provide flexible options to meet the diverse needs of the ML community. MG Analyzer is built on Turtle language and Notation3 reasoning engine, envisioned as a first step toward Linked Open Data for ML workflow management. We have also encoded our proposed model licenses into rules and demonstrated the effects of GPL and other commonly used licenses in model publishing, along with the flexibility advantages of our licenses, through comparisons and experiments.
AI Licensing, Semantic Web, Open Source, Responsible AI
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792968ccs: Social and professional topics Testing, certification and licensingccs: Software and its engineering Open source modelccs: Theory of computation Automated reasoning
Resource Availability:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.18277024.
@@ 1. Introduction
In recent years, the compelling generalization capabilities provided by billion-parameter models (Wei et al., 2023), along with the high computational and data costs associated with their training (Maslej et al., 2024), have motivated ML developers to collaborate incrementally rather than train models from scratch. A common approach is to download a Pre-Trained Model (PTM) (Jiang et al., 2023) and fine-tune it for downstream task (Hu et al., 2022). However, these practices can pose legal risks if the use or redistribution of reused components violates their licenses, akin to GPL violations in Open Source Software (OSS) (Mathur et al., 2012). Failure to comply with license terms can result in termination and legal action [^1] See CoKinetic Systems, Corp. v. Panasonic Avionics Corporation, 1:17-cv-01527.. Another concern is the license chosen for republishing the work. Some developers adhere to traditional software publishing practices and select OSS licenses for their models (Devlin et al., 2019; Ni et al., 2022; Duan et al., 2025), which often lack clear definitions and conditions regarding ML activities and do not effectively prevent undesirable use. For example, a licensee can close-source your published models, even if they are licensed under GPL, without violating any terms.
@@
There are three possible ways to mitigate above risks. First, developers could avoid using any third-party materials. However, this is extremely difficult for individual developers, as training PTMs is expensive and requires vast amounts of data. For instance, the training dataset for GPT-2 (Radford et al., 2019) was collected from 45 million web pages, governed by various licenses and terms of use. Second, a new publishing standard for ML projects could be developed, which might include drafting specific licenses for models and datasets (Benjamin et al., 2019; Contractor et al., 2022), along with a compatibility table to guide their reuse policies. However, this approach does little to address existing conflicts in ML projects that already rely on assets released under traditional licenses. Furthermore, it is impractical to expect all publishers to relicense their previous works. Third, we can scan the reused assets in ML projects and analyze existing license noncompliance to eliminate them. This is a common solution applied to OSS projects (Jaeger et al., 2017) but it cannot be directly extended to ML projects, as they can involve complex coupling mechanisms and different licensing frameworks that are interwoven within an ML development workflow.
@@
Take MixLoRA (Li et al., 2024) as an example: it is licensed under Apache-2.0 (an OSS license) and is fine-tuned on Llama 2 model (Touvron et al., 2023) (governed by Llama 2 Community License (Meta Platforms, 2024), a model license) using the Cleaned Alpaca Dataset (Taori et al., 2023), which is licensed under CC BY-NC-4.0 (a free-content license from Creative Commons, a.k.a CC (Commons, 2024)). Previous OSS license analysis tools (Mathur et al., 2012; Ombredanne, 2020) that only consider package reference dependencies and focus on software licenses will fall short in such ML scenarios. Therefore, to provide license analysis for ML projects, the key is to develop a specific ontology that describes the ML workflow and provides a corresponding interpretative solution for licenses, covering all licensing frameworks and disambiguating their mapping rules related to ML activities. Moreover, the lack of consensus in standard model publishing practices and the inflexibility of existing publicly available model licenses have led many developers to publish their models under OSS licenses or even free-content licenses (Groeneveld et al., 2024; CohereForAI, 2024), further complicating the design of license analysis methods.
@@
In this paper, we propose a two-pronged approach to address these challenges. First, to resolve existing noncompliance in ML projects, we introduce MG Analyzer, a tool that constructs ML workflows as Resource Description Framework (RDF) (Pan, 2009) graphs and assesses potential license compliance issues, improper license selection, granting of rights, restrictions, and obligations related to ML asset reuse. Second, we propose MG Licenses, a new set of model-specific licenses with CC-style licensing options, to promote standardized model publishing. To present potential risks of using OSS, model, and free-content licenses in model publishing scenarios, we evaluate them with the MG Analyzer on a typical workflow. We also demonstrate the flexibility of MG Licenses in encompassing nearly all licensing conditions provided by other model licenses through comparisons. The main contributions of our paper are:
@@
  • We identified the challenges of license compliance and model licensing in ML workflow management.
@@
  • We developed MG Analyzer to automate license analysis in ML projects. It includes a vocabulary for describing ML workflows with dependencies and license rules, and provides an interface to convert user-input workflow descriptions into RDF graphs. These graphs enable the tool to construct dependencies, reason, and detect license conflicts.
@@ 2. Background and Related Work
2.1. ML Workflow Management
Accurate description and management of the provenance and dependencies of ML assets are critical for legal compliance analysis in ML workflows (Wang et al., 2025). Currently, there are two active directions toward standardized ML workflow management. The first is the development of AI Bills of Materials (BOM) formal description frameworks designed to enumerate ML assets, parameters, and procedures required to develop AI models. Among these efforts, Software Package Data Exchange 3.0 (SPDX 3.0) stands out as an open standard for expressing AI BOMs (Bennet et al., 2024). SPDX 3.0 provides a graph-based, extensible data model that enables detailed provenance tracking and license metadata across complex software and AI development pipelines. CycloneDX [^2] CycloneDX Bill of Materials Standard is another emerging standard focused on cybersecurity and supply chain risk, supporting metadata for vulnerability disclosure, threat modeling, and policy enforcement. While these standards offer a strong foundation for traceability and transparency, they remain general-purpose and lack native support for domain-specific semantics required in ML legal compliance analysis. In particular, they do not provide built-in vocabulary for ML-specific development actions such as distillation, fine-tuning, or embedding, nor do they capture legal concepts such as granted rights, reserved rights, and license conditions in a machine-interpretable form.
The second is the development of MLOps toolkits, such as ModelDB (Vartak and Madden, 2018), MLflow (Zaharia et al., 2018), Texera (Wang et al., 2024b; Huang et al., 2024), Neptune333neptune.ai, and SAS Model Manager [^4] SAS Model Manager. These platforms facilitate the end-to-end management of ML workflows by supporting functionalities such as experiment tracking, model versioning, performance monitoring, and deployment orchestration. While they enhance reproducibility, governance, and automation in ML lifecycle management, they typically lack native support for legal metadata tracking or license compliance verification, requiring integration with external standards like SPDX 3.0 for comprehensive workflow analysis. In parallel, recent knowledge graph-based efforts such as CRUX (Wang et al., 2022) and ModsNet (Wang et al., 2023, 2024a) construct ML workflow knowledge bases to support workflow-centric queries, such as retrieving the best-performing models for a given dataset. Other efforts, such as AiiDA (Huber et al., 2020) and ProvONE (Cao et al., 2016), focus on fine-grained provenance tracking in scientific workflows. They provide graph-based models to capture process and data lineage. Inspired by these approaches, our proposed ML workflow representation adopts their principles but extends them to support automated license compliance analysis by formally representing the reasoning logic of license rules and analyzing dependencies among ML assets.
2.2. ML Model Licensing
@@ 3. MG Analyzer
Figure 1. Overview of MG Analyzer (”mg” is the prefix for our proposed vocabulary).
This section aims to introduce the specific design of MG Analyzer by exploring three questions: (i) How should we represent the workflows of ML projects to facilitate license-related noncompliance analysis? (ii) How do we establish a mapping from license text to reasonable rules? (iii) What types of license compliance issues can arise in ML projects, and how can we detect them? Before delving into the detailed design that answers these questions, we first provide an overview that serves as a roadmap for this section. As illustrated in Figure 1, the process of MG Analyzer is divided into three main parts: Construction, Reasoning, and Analysis.
@@ 3.1. ML Workflow Representation
①ML workflows often involve various components (e.g., code, datasets, images, model weights, services), each governed by licenses from different frameworks. Additionally, non-standard licensing practices are prevalent in current ML projects (Rajbahadur et al., 2021; Duan et al., 2024), for instance, C4AI Command R+ model (CohereForAI, 2024) is licensed under a free-content license: CC-BY-NC-4.0. Therefore, the representation should be flexible enough to cover such situations.
@@
At this stage, we can describe a base ML workflow, as illustrated [^5] For clarity, the “mg” prefix is omitted, and certain properties are merged or filtered. schematically in Figure 2 (The RDF graph can be referred to in gen.ttl in Figure 1). In this base workflow, the derived input and output of each Action can be represented as blank nodes, serving as placeholders. These placeholders will be populated by reasoning the output yielded by the previous action and the input for the current action, respectively.
@@
The rights-using dependencies describe the rights that must be granted for actions performed on works. For example, when executing a Train action on a model, it requires the rights to use and modify (termed as Usage in MG Vocabulary) from the model’s license [^6] An example of such rights granting can be found in the Llama 2 Community License, which states, ”You are granted a … to use, reproduce, distribute, copy, create derivative works of, and make modifications to the Llama Materials.”.. Similarly, the rights-using dependencies should proliferate according to compositional dependencies. For instance, the requirement to modify a model extends to all its submodels. In the MG Analyzer, we create new nodes called Request to represent this dependency. Additionally, it is insufficient to only check the granting rights; the reserved rights must also be verified. Depending on the clarity of the license text, some rights may either be explicitly granted or reserved. Furthermore, certain license clauses can waive the requirement for specific rights. For instance, both GPL-3.0 and CC licenses include automatic relicensing clauses for downstream recipients, which eliminate the need for a sublicense right.
@@ 3.2. License Rule Encoding and Reasoning
The final aspect is applicable term generalization, where we encode the triggering conditions of a license term into the following properties: range of input work forms, range of output work forms, and types of actions. Figure 3 illustrates an example of the GPL-3.0 derivatives rule [^7] The original GPL-3.0 license text reads: ”You may convey a work based on the Program … in the form of source code … provided that you also meet all of these conditions: … stating that you modified it … it is released under this License … keep intact all notices … license the entire work, as a whole, under this License …” ., which represents the Combine action applied to input works in code format, resulting in a code output. This action triggers the ”derivative” clause in GPL-3.0, indicating that the license of the output work must be compatible with GPL-3.0 (e.g., AGPL-3.0). Additionally, five restrictions must apply to the output work if it is to be republished, as dictated by this definition dependency. This rule does not include any Use Restrictions, which apply to output works regardless of whether they are republished. We found this subtle distinction to be crucial in ML license analysis, as most OSS license terms are triggered by distribution, and their definitions of ”distribution” typically exclude publishing as a service. However, in the case of models, the common deployment method is through a web interface, leading to many OSS license restrictions being circumvented in such scenarios (ref Section 5).
Furthermore, multiple rules may lead to the same output definition, and we provide the option of fuzz form matching by enabling fuzz rules to enhance the interpretive capabilities of the MG Analyzer. It is worth mentioning that the reasoning results behind the restrictions derived from these rules require further analysis for validation. Taking the rule in Figure 3 as an example, the presence of additional discrimination terms in the final work that violate GNU freedom (Greenbaum, 2016) dictates whether errors related to GNU freedom conflicts should be reported. We present snippets of our encoded rules written in Turtle format in Appendix B. The list of supported licenses, whose terms have been encoded in MG Analyzer, is shown in Table 2, covering nearly all top-ranking licenses for published models on HuggingFace [^8] https://huggingface.co/models, a popular model publishing platform.
@@
With MG Vocabulary, the ML workflow with compositional dependencies, and encoded license rules, we can reason to derive the definition and rights-using dependencies, thereby determining the applicable licenses for intermediate works. The license determination in the MG Analyzer follows an incremental and minimal noncompliance strategy, where the new license only applies to the incremental parts of the work without affecting the original work (a common practice in licensing). Furthermore, to avoid introducing additional compliance issues during analysis, we use Unlicense as the default license when applicable. However, an exception arises with the license proliferation clauses found in copyleft licenses, such as GPL-3.0 and OSL-3.0, which require that the entire new work be licensed under the same terms. Our analyzer incorporates reasoning logic to identify applicable licensing solutions (a snippet of logic can be found in Appendix B), but unresolved conflicts may occur if multiple copyleft clauses are triggered. In such cases, the MG Analyzer will select one of these copyleft licenses and report an error during the Analysis Stage.